# CATP: Contextually Adaptive Token Pruning for Efficient and Enhanced Multimodal In‑Context Learning

### v1

@@ Introduction
Benchmarks. To validate CATP, we conduct extensive experiments on eight vision-language benchmarks. They include standard multimodal ICL evaluations (Li 2025): VQAv2 (Goyal et al. 2017), VizWiz (Gurari et al. 2018), OK-VQA (Marino et al. 2019), GQA (Hudson and Manning 2019), and Hatefulmemes (Kiela et al. 2020). Three latest and challenging multimodal benchmarks are also used: MMStar (Chen et al. 2024a), MME‑Realworld (Zhang et al. 2024b), and VL‑ICL (Zong, Bohdal, and Hospedales 2025).
Baselines and models. We compare CATP with the vanilla model, random pruning, and seven strong baselines. In addition to the previously introduced FastV, FitPrune, VTW, and DivPrune, we also evaluate HiRED (Arif et al. 2025), SparseVLM (Zhang et al. 2024a), and PLPHP (Meng et al. 2025). For methods relying on user‑specified hyperparameters, we test four configurations and report the best performance. We first evaluate each pruning method on LLaVA‑Next‑7B under a range of pruning ratios. We then report additional results at a fixed ratio on LLaVA‑Next‑13B, InternVL2.5‑8B, and Qwen2.5VL‑7B. In LLaVA-Next every image in the input sequence is uniformly converted into 576 tokens, whereas in InternVL2.5-8B and Qwen2.5VL-7B, the number of tokens per image depends on its original resolution and ranges from 576 to 1280. Appendix 3.1 provides further introductions to the benchmarks, baselines, models, and associated processing details.
Implementation details. All experiments are conducted on 4 NVIDIA H200 GPUs. For every benchmark the validation samples serve as query samples. Each query is paired with four ICDs randomly retrieved from the training split, producing a 4-shot sequence. For LLaVA-Next-7B, InternVL2.5-8B, and Qwen2.5VL-7B, we set $K=6$. For LLaVA-Next-13B, we set $K=10$. Across all LVLMs we set $\lambda_{1}=0.7$ and $\lambda_{2}=0.6$. As CATP operates without the full attention matrices, it is compatible with FlashAttention (Dao et al. 2022), which further boosts its efficiency.
@@
Table 1 reports the performance of CATP under three commonly used pruning ratios. These are the levels at which pruning methods typically bring real efficiency gains. CATP achieves the best performance under all three settings. At both 66.7% and 77.8%, it enhances performance over the vanilla model. When the ratio reaches 89.9%, CATP incurs only a modest 1.4% drop. Meanwhile, CATP is effective across diverse LVLMs and benchmarks, as shown in Tables 1 and 2. On four LVLMs and eight benchmarks, CATP consistently achieves the best performance. Notably, none of the seven SOTA baselines improve multimodal ICL, with some even performing worse than random pruning. In contrast, CATP achieves performance enhancement by capturing complex interactions and identifying the most important image tokens for the overall ICL process.
@@
To demonstrate the efficiency of CATP, we conduct extensive comparative experiments on four LVLMs, reporting TFLOPs, inference latency, and KV Cache size. Following FastV, we compute the FLOPs of the multi-head attention and feedforward network modules as $4nd^{2}+2n^{2}d+3ndm$, where $n$ is the number of image tokens, $d$ is the hidden state size and $m$ is the intermediate size of the FFN. For latency, we report the average total inference time per sequence. For KV Cache memory usage, we report the average GPU memory consumption during inference. Table 3 reports the results for each LVLM, averaged across eight benchmarks. CATP achieves the highest accuracy and shows excellent efficiency. It ranks second only to DiVPrune in reducing FLOPs and KV Cache and delivers the largest cut in inference latency. As pointed out by (Wen et al. 2025a), latency is the most reliable indicator of token pruning efficiency because the execution cost of the method is not reflected in FLOPs or KV Cache. These results indicate that CATP brings real efficiency gains in multimodal ICL. Together with its unique performance boost, CATP is an outstanding pruning method with strong practical potential.
@@
Adaptability to diverse in-context sequence configurations. In practical multimodal ICL, both the number of ICD shots and the ICD selection strategy vary across applications. Since sequence construction strongly affects performance (Gao, Fisch, and Chen 2021; Liu et al. 2021), we evaluate CATP under diverse configurations to assess its robustness. Figure 5(a) shows that CATP performs better as shot count increases, due to its ability to handle rising token redundancy and enhance reasoning by keeping only informative tokens. As LVLMs support longer contexts, CATP’s benefits can grow accordingly. Its inference speedup also scales nonlinearly with shot count (see Appendix 3.2). Figure 5(b) reports results under four ICD selection strategies. RS denotes the random sampling used in our main experiments. I2I and IT2IT are similarity-based methods commonly used in modern retrieval-augmented generation (RAG) systems, based on image-only and joint image-text similarity, respectively. TACO (Li et al. 2025c) is a SOTA ICD selection model. CATP brings greater performance gains as the quality of the sequence improves, suggesting that it can be integrated into existing systems to provide incremental benefits.
@@ Conclusion
In this paper, we introduce CATP, a novel and training-free image token pruning method tailored to multimodal ICL. In the two-stage pruning process, CATP adaptively identifies the image tokens that are most important to the entire ICL process based on the input in-context sequence, and discards the rest. This leads to improvements in both performance and efficiency. CATP fills the gap in existing work on multimodal ICL and offers a reliable and efficient solution to enhance the capability of LVLMs when handling complex inputs. We believe CATP provides a solid foundation and valuable insights for advancing LVLM capabilities.
@@ 2   Attention Shifts in Multimodal ICL
Attention shift (Cao et al. 2022; Zhang et al. 2025b) is the layer-wise drift of cross-modal attention within LVLM’s decoder: visual tokens that initially draw little attention can gradually absorb a large fraction of the text-to-vision attention budget, so the amount of attention they receive no longer reflects their true saliency. Because transformers renormalize attention at every block, even small relocations compound, so by the final layers the pattern of “important” image tokens can be almost inverted relative to the shallow view. Empirically, this drift appears as clusters of redundant patches and tends to concentrate attention weights disproportionately on tokens that are closer to the text, whereas distinctive but more distant regions are overlooked, as illustrated in Figure 9.
@@
Metrics that rely on raw attention therefore inherit this instability. Average or cumulative attention scores, top-k frequency counts, and similar heuristics implicitly assume that a token’s weight is stationary. When that assumption fails, the metrics overestimate tokens that merely inherit migrated attention and underestimate those whose relevance surfaces late. Pruning or saliency mapping driven by such metrics either retains many near-duplicate patches, wasting computation, or excises genuinely informative ones, degrading vision-language performance. The mismatch between predicted and actual redundancy also grows with model depth.
The problem worsens in multimodal ICL where images and text are interleaved. Relative-position or rotary encodings bias attention toward tokens that sit closest to the current textual span, so each newly inserted image block initially receives an artificial boost. As more text and images accumulate, attention repeatedly migrates toward the most adjacent visual tokens, pushing earlier images into a low-attention regime regardless of their semantic relevance. This positional bias amplifies attention shifts, making importance estimates increasingly local and myopic, and causing pruning methods based on attention to fail far more often in long, interleaved contexts than in single-image tasks.
@@ 3.1 Setup Details
  • FastV: FastV shows that in the middle and deep layers of LVLMs, image tokens receive much less attention than text tokens, which indicates severe redundancy in these layers. The method therefore concentrates on early-stage token pruning by exploiting the attention that image tokens obtain from all other tokens. We apply FastV at the 2nd layer of the decoder.
@@ 3.2 Efficiency with More Shots
As shown in Table 5, the efficiency gains brought by CATP become more evident as the ICD shot count in the input sequence increases. With the growing demand for long-context tasks, the context length of current generation LVLMs has greatly expanded. This extension permits the inclusion of many multimodal ICDs, more fully capturing user needs and even enabling many-shot ICL, which is an important part of retrieval-augmented generation (RAG) systems. CATP can be effectively incorporated into these scenarios to mitigate inference latency, thereby promoting its adoption in a broader range of industrial-grade applications.
@@ 3.4 Analysis of $\lambda_{1}$ and $\lambda_{2}$
When $\lambda_{1}$ increases, Stage 1 of CATP places greater emphasis on image–text alignment. When $\lambda_{2}$ increases, Stage 2 prunes the image tokens in the ICDs with heightened focus on semantic relevance. Figure 11 provides a closer look at how the two balancing coefficients affect different LVLMs. Panel (a) shows that a moderate emphasis on image-text alignment ($\lambda_{1}=0.7$) yields the largest relative gain of about two percentage points for InternVL2.5-8B, whereas LLaVA-Next-7B prefers either a lower or the maximal setting and Qwen2.5VL-7B benefits most when $\lambda_{1}$ approaches 0.9. Panel (b) indicates that giving more weight to semantic relevance in Stage 2 brings clear improvements only up to a point. InternVL2.5-8B again peaks at $\lambda_{2}=0.7$; further increases gradually erode its advantage. LLaVA-Next-7B reaches its best value near $\lambda_{2}=0.9$, while Qwen2.5VL-7B performs best at the smallest $\lambda_{2}$ and declines steadily afterwards. These trends confirm that the relative importance of alignment and relevance signals is model‐specific.

### v2

@@ Introduction
Benchmarks. To validate CATP, we conduct extensive experiments on eight vision-language benchmarks. They include standard multimodal ICL evaluations (icl3): VQAv2 (vqav2), VizWiz (VizWiz), OK-VQA (OK-VQA), GQA (GQA), and HatefulMemes (hateful). Three latest and challenging multimodal benchmarks are also used: MMStar (mmstar), MME‑Realworld (mme), and VL‑ICL (vlicl).
Baselines and models. We compare CATP with the vanilla model, random pruning, and seven strong baselines. In addition to the previously introduced FastV, FitPrune, VTW, and DivPrune, we also evaluate HiRED (att2), SparseVLM (att1), and PLPHP (plphp). For methods relying on user‑specified hyperparameters, we test four configurations and report the best performance. We first evaluate each pruning method on LLaVA‑Next‑7B under a range of pruning ratios. We then report additional results at a fixed ratio on LLaVA‑Next‑13B, InternVL2.5‑8B, and Qwen2.5VL‑7B. In LLaVA-Next every image in the input sequence is uniformly converted into 576 tokens, whereas in InternVL2.5-8B and Qwen2.5VL-7B, the number of tokens per image depends on its original resolution and ranges from 576 to 1280. Appendix 3.1 provides further introductions to the benchmarks, baselines, models, and associated processing details.
Implementation details. All experiments are conducted on 4 NVIDIA H200 GPUs. For every benchmark the validation samples serve as query samples. Each query is paired with four ICDs randomly retrieved from the training split, producing a 4-shot sequence. For LLaVA-Next-7B, InternVL2.5-8B, and Qwen2.5VL-7B, we set $K=6$. For LLaVA-Next-13B, we set $K=10$. Across all LVLMs, we set $\lambda_{1}=0.7$ and $\lambda_{2}=0.6$. As CATP operates without the full attention matrices, it is compatible with FlashAttention (flash), which further boosts its efficiency.
@@
Table 1 reports the performance of CATP under three commonly used pruning ratios. These are the levels at which pruning methods typically bring real efficiency gains. CATP achieves the best performance under all three settings. At both 66.7% and 77.8%, it enhances performance over the vanilla model. When the ratio reaches 89.9%, CATP incurs only a modest 1.4% drop. Meanwhile, CATP is effective across diverse LVLMs and benchmarks, as shown in Tables 1 and 2. On four LVLMs and eight benchmarks, CATP consistently achieves the best performance. Notably, none of the seven SOTA baselines improve multimodal ICL, with some even performing worse than random pruning. In contrast, CATP achieves performance enhancement by capturing complex interactions and identifying the most important image tokens for the overall multimodal ICL process.
@@
To demonstrate the efficiency of CATP, we conduct extensive comparative experiments on four LVLMs, reporting TFLOPs, inference latency, and KV Cache size. Following FastV, we compute the FLOPs of the multi-head attention and feedforward network modules as $4nd^{2}+2n^{2}d+3ndm$, where $n$ is the number of image tokens, $d$ is the hidden state size and $m$ is the intermediate size of the FFN. For latency, we report the average total inference time per sequence. For KV Cache memory usage, we report the average GPU memory consumption during inference. Table 3 reports the results for each LVLM, averaged across eight benchmarks. CATP achieves the highest accuracy and shows excellent efficiency. It ranks second only to DivPrune in reducing FLOPs and KV Cache and delivers the largest cut in inference latency. As pointed out by (pruning2), latency is the most reliable indicator of token pruning efficiency because the execution cost of the method is not reflected in FLOPs or KV Cache. These results indicate that CATP brings real efficiency gains in multimodal ICL. Considering both its unique performance boost and its substantial efficiency gains, CATP stands out as an effective pruning method with strong practical potential for interleaved image-text tasks.
@@
Adaptability to diverse in-context sequence configurations. In practical multimodal ICL, both the number of ICD shots and the ICD selection strategy vary across applications. Since sequence construction strongly affects performance (sen2; sen1), we evaluate CATP under diverse configurations to assess its robustness. Figure 5(a) shows that CATP performs better as shot count increases, due to its ability to handle rising token redundancy and enhance reasoning by keeping only informative tokens. As LVLMs support longer contexts, CATP’s benefits can grow accordingly. Its inference speedup also scales nonlinearly with shot count (see Appendix 3.2). Figure 5(b) reports results under four ICD selection strategies. RS denotes the random sampling used in our main experiments. I2I and IT2IT are similarity-based methods commonly used in modern retrieval-augmented generation (RAG) systems, based on image-only and joint image-text similarity, respectively. TACO (taco) is a SOTA ICD selection model. CATP brings greater performance gains as the quality of the sequence improves, suggesting that it can be integrated into existing systems to provide incremental benefits.
@@ Conclusion
In this paper, we introduce Contextually Adaptive Token Pruning (CATP), a novel and training-free image token pruning method tailored to multimodal ICL. In the two-stage pruning process, CATP adaptively identifies the image tokens that are most important to the entire ICL process based on the input in-context sequence and discards the rest. This leads to improvements in both performance and efficiency. CATP addresses an important gap in existing work on multimodal ICL and offers a reliable and efficient solution to enhance the capability of LVLMs when handling complex inputs. We believe CATP provides a solid foundation and valuable insights for advancing LVLM capabilities.
Acknowledgments
We acknowledge the computing resources provided by NSF ACCESS.
Figure 8: Overview pipeline of CATP. CATP consists of two stages. Stage 1 is applied between the projector and the decoder. It measures the importance of image tokens using feature distribution diversity and alignment with their paired text. Stage 2 operates in two shallow decoder layers. It first prunes the ICD image tokens by combining inter-layer attention differences with semantic relevance, and then prunes the query sample’s image tokens in the following layer.
@@ 2   Attention Shifts in Multimodal ICL
Attention shift (as2; limi1) is the layer-wise drift of cross-modal attention within LVLM’s decoder: visual tokens that initially draw little attention can gradually absorb a large fraction of the text-to-vision attention budget, so the amount of attention they receive no longer reflects their true saliency. Because transformers renormalize attention at every block, even small relocations compound, so by the final layers the pattern of “important” image tokens can be almost inverted relative to the shallow view. Empirically, this drift appears as clusters of redundant patches and tends to concentrate attention weights disproportionately on tokens that are closer to the text, whereas distinctive but more distant regions are overlooked, as illustrated in Figure 9.
@@
Thus, metrics that rely on raw attention inherit this instability (ada; logic). Average or cumulative attention scores, top-k frequency counts, and similar heuristics implicitly assume that a token’s weight is stationary. When that assumption fails, the metrics overestimate tokens that merely inherit migrated attention and underestimate those whose relevance surfaces late. Pruning or saliency mapping driven by such metrics either retains many near-duplicate patches, wasting computation, or excises genuinely informative ones, degrading vision-language performance. The mismatch between predicted and actual redundancy also grows with model depth.
The problem worsens in multimodal ICL, where images and text are interleaved. Relative-position or rotary encodings bias attention toward tokens that sit closest to the current textual span, so each newly inserted image block initially receives an artificial boost. As more text and images accumulate, attention repeatedly migrates toward the most adjacent visual tokens, pushing earlier images into a low-attention regime regardless of their semantic relevance. This positional bias amplifies attention shifts, making importance estimates increasingly local and myopic, and causing pruning methods based on attention to fail far more often in long, interleaved contexts than in single-image tasks.
@@ 3.1 Setup Details
  • FastV: FastV shows that in the middle and deep layers of LVLMs, image tokens receive much less attention than text tokens, which indicates severe redundancy in these layers. Therefore, the method concentrates on early-stage token pruning by exploiting the attention that image tokens obtain from all other tokens. We apply FastV at the 2nd layer of the decoder.
@@ 3.2 Efficiency with More Shots
As shown in Table 5, the efficiency gains brought by CATP become more evident as the ICD shot count in the input sequence increases. With the growing demand for long-context tasks, the context length of current generation LVLMs has greatly expanded. This extension permits the inclusion of many multimodal ICDs, more fully capturing user needs and even enabling many-shot ICL, which is an important part of retrieval-augmented generation (RAG) systems. CATP can be effectively incorporated into these scenarios to mitigate inference latency, thereby promoting its adoption in a broader range of industrial-grade applications (xiong1; xiong2).
@@ 3.4 Analysis of $\lambda_{1}$ and $\lambda_{2}$
When $\lambda_{1}$ increases, Stage 1 of CATP places greater emphasis on image-text alignment. When $\lambda_{2}$ increases, Stage 2 prunes the image tokens in the ICDs with heightened focus on semantic relevance. Figure 11 provides a closer look at how the two balancing coefficients affect different LVLMs. Panel (a) shows that a moderate emphasis on image-text alignment ($\lambda_{1}=0.7$) yields the largest relative gain of about two percentage points for InternVL2.5-8B, whereas LLaVA-Next-7B prefers either a lower or the maximal setting and Qwen2.5VL-7B benefits most when $\lambda_{1}$ approaches 0.9. Panel (b) indicates that giving more weight to semantic relevance in Stage 2 brings clear improvements only up to a point. InternVL2.5-8B again peaks at $\lambda_{2}=0.7$; further increases gradually erode its advantage. LLaVA-Next-7B reaches its best value near $\lambda_{2}=0.9$, while Qwen2.5VL-7B performs best at the smallest $\lambda_{2}$ and declines steadily afterwards. These trends confirm that the relative importance of alignment and relevance signals is model‐specific.